OpenStudy (anonymous):

solve the equation for y , then find the value of y for each value of x.. 1) 5x=-4y+4 ; x=1,2,3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[5x=-4y+4\\ 5x-4=-4y\\ -\frac{5}{4}x+1=y\]

OpenStudy (misty1212):

steps were a) subtract 4 b) divide by -4

OpenStudy (anonymous):

THANK YOU ;)